OpenStudy (anonymous):

This is for my little bro. Solve the inequality. Explain all steps and identify the properties used. 3a−4≤5

OpenStudy (midhun.madhu1987):

Add 4 to both sides of the equation. Then divide both sides by 3 and u get the value of a.

OpenStudy (carson889):

Essentially you want to separate the variable (3a) from the constants ( -4 and 5). You can treat the inequality similar to an equal sign, in that, what you do one side, you must do to the other. Therefore, the first step to separate the constants from the variable is to add 4 to both sides: 3a - 4 + 4 <= 5 + 4 Now it looks like this: 3a <= 9 Now we want to isolate the variable ( basically get rid of the multiple three on the left side of the inequality). To do this we simply divide both sides by 3: 3a/3 <= 9/3 a <= 3 There you have it, you've solved the inequality. Three is greater than or equal to a.